what is the quadratic equation of the co-orinates: Turning point= (7,3) and x-intercepts= (6,8)
Answer by khwang(438)

Since the turning point (vertex) is (7,3), assume the quadratic equation
as: y - 3 = k(x-7)^2.
Next, since it passes through (6,8), we have
8 - 3 = k(6-7)^2 = k, or k = 5.
The equation is y - 3 = 5(x-7)^2.
Kenny